OpenStudy (anonymous):

The scale on this map is 2 cm for 300 m. Using a ruler, the distance from Dallas Public Library to City Hall measures 1.8 cm. What is the approximate distance of these two locations?

jimthompson5910 (jim_thompson5910):

2 cm/300 m = 1.8 cm/ x m 2/300 = 1.8/x 2x = 300*1.8 2x = 540 x = ??